A STANDARD FORM FOR INCOMPRESSIBLE SURFACES IN A HANDLEBODY

LINDA E. GREEN

Abstract. Let Σ be a compact surface and let I be the unit interval. This paper gives a standard form for all 2-sided incompressible surfaces in the 3-manifold Σ×I. Since Σ × I is a handlebody when Σ has boundary, this standard form applies to incompressible surfaces in a handlebody.